# references.bib

### v3

@@ 4 Results
We describe our findings for the three models introduced, on the Mnist, FashionMnist, Svhn and Cifar10 datasets. In particular, we focus on the unique properties of representations obtained within the FF model, i.e. a model trained with Forward-Forward on its natural goodness objective. Such properties – e.g. the emergence of category-specific ensembles and the presence of shared units across them – establish a link between neural networks trained with the Forward-Forward algorithm and biological cortical networks described in subsection 2.2.
@@ 4.1 Classification accuracy
Before delving into the main results of this work, we evaluate the performances of our models on the classification tasks at hand. Table 1 contains results in terms of test set classification accuracy for all models we employed – FF, BP/FF and BP – on Mnist, FashionMnist, Svhn and Cifar10. While some of these accuracy values are far from the state-of-the-art (i.e., respectively, $0.997$ [ciresan2010bigdeep], $0.931$ [xiao2017fashion], $0.860$ [pitsios2017] and approximately $0.7$ [lin2015far], for fully-connected networks), they are a solid ground on which to build our subsequent investigations. Training details and hyperparameters for all models are reported in Appendix C.
@@ 4.2 Forward-Forward elicits sparse neuronal ensembles
Figure 2: Sparsity of category-specific representations. We report the sparsity of representations - computed as described in subsection 3.5 - for the three models FF, BP/FF and BP on the Mnist dataset (same models as in the first row of Table 1). Sparsity values are the average on 10 runs.
The FF and BP/FF models – based on the original Forward-Forward network architecture, and trained according to the goodness objective (subsection 3.2 and subsection 3.3) – exhibit typically high sparsity levels in their representations, in clear contrast with BP (see Figure 2 and Table 2).
Table 2: Average sparsity for all combinations of model, dataset and layer, according to the definition given in subsection 3.5. Results are expressed as mean $\pm$ std. dev. computed over 10 runs with independent random weights initialization.
@@
When the sparsity level is sufficiently high (S > 0.5) we are able to identify small sets of neurons (ensembles) that consistently co-activate across all the samples of the same class, similarly to what has been observed in cortical representations [yuste2015neuron, miller2014visual, harris2005neural].
Figure 1 (Panels B, C) shows an example of average neuron activations for each class in Layer 1 of the FF model trained on Mnist, and showcases the emergence of sparse, category-specific, ensembles (see the Appendix D for a similar visualisation for Layers 2 and 3 of the same model and Appendix E for Layer 1 in all the models).
These representations often elicit a remarkably low fraction of active units: ensembles composed by a few percent of the neurons in a layer are typically observed both for the simplest datasets (Mnist, FashionMnist) and for more complex datasets (Svhn, Cifar10), with a slight tendency of the FF model to create larger ensembles in the latter case (Table 3).
@@
4.3 Semantically similar classes elicit ensembles with shared neurons
Drawing a parallel with a phenomenon observed in Neuroscience [yoshida2020natural], related categories can be expected to share units of their ensembles. This is indeed what we observe, as shown in Figure 3. Results are reported for FashionMnist, where different classes of clothes or shoes may contain a common share of visual features. In this regard, we observe a clear tendency to share units between similar classes – e.g. across representations of pullover, coat and shirt.
In the following section, we provide evidence that a unit can be shared across two ensembles even if one of these refers to an unseen category (i.e., excluded from the training set but whose representation, extracted at test time, generates a valid ensemble), as we show in Figure 4 (Panel C).
Figure 3: Semantically similar classes in FashionMnist can elicit ensembles with shared neurons.
@@ 4.3 Semantically similar classes elicit ensembles with shared neurons
Panel B Element $i,j$ of the matrix indicates how many units are shared between the ensembles of category $i$ and category $j$ (normalized by the ensemble sizes), by using the Jaccard similarity index: $J(\mathcal{E}^{i},\mathcal{E}^{j})=\frac{\mid\mathcal{E}^{i}\cap\mathcal{E}^{j}\mid}{\mid\mathcal{E}^{i}\cup\mathcal{E}^{j}\mid}$. The results are referred to a single training run.
@@ 4.4 Representations of unseen categories can elicit well-defined ensembles
We investigate the ability of a trained FF model to respond to unseen categories with a coherent activation pattern which is typical of the ensembles we found on the categories seen at training time. To this end, we repeatedly train FF on FashionMnist, removing one category at a time. Then, we extract the representation of the missing category, and verify if a valid ensemble is formed. We find that in all the ten cases, a new ensemble is formed, sharing the same characteristics of the ones emerging for seen categories, apparently with the only exception of a lower average activation of their constituent units (see Figure 4 for one example, and the Appendix F for a more detailed account).
In several cases, we also found that the ensembles of unseen categories share units with the ensembles of seen categories, when endowed with similar visual features or semantics (Figure 4, Panel C). A more extensive exploration of these cases is also reported in the Appendix F.
Figure 4: The representation of unseen categories forms a valid ensemble in FF trained on FashionMnist.
@@
5 Discussion and conclusions
The main finding of our work is that artificial neural networks – trained with the Forward-Forward algorithm – can elicit biologically plausible representations in the form of sparse neuronal ensembles [yuste2015neuron, miller2014visual], which robustly and reliably encode meaningful information. Such ensembles can also share units across semantically related inputs [yoshida2020natural].
We started our investigation by collecting and analysing representations from FF networks trained on Mnist, FashionMnist, Svhn and Cifar10, and defined, separately for each category, subsets of units (ensembles) that prominently and consistently activate in response to data in such category (subsection 4.2). These category-specific ensembles turn out to be composed of a few units, which is consistent with experimental findings on the sensory cortex [miller2014visual]. Furthermore, when image categories are characterised by a certain degree of visual similarity, the corresponding ensembles often share one or more units (Figure 3, Panel B). These results are consistent with phenomena observed in sensory cortices [yoshida2020natural] – as discussed in detail in subsection 2.2 – where ensembles are defined.
We then tested the ability of trained FF models to cope with new data, and we observed that the activations in response to an unseen input category formed in many cases a new ensemble, with similar characteristics of sparsity as the ones formed for other classes during training (Figure 4, Panels A and B). Furthermore, we noticed that the ensembles of unseen categories can share units with the ensembles of semantically-related categories used as training data (Figure 4, Panel C). These findings suggests that FF can perform well in zero-shot classification tasks, which is particularly relevant in view of the importance of zero/few-shot learning in human and animal cognitive performances [lake2015human].
While absent in the BP model, the existence of ensembles composed by a few units is not unique to FF. It is indeed observed also in BP/FF (subsection 4.2), where the ensembles turns out to be of comparable size. However, despite their similarity, the FF and BP/FF models are profoundly different, as demonstrated by the different interplay between inhibition and excitation in these models (see Appendix, G). We observe in this regard that the excitatory/inhibitory (E/I) balance plays a key role in the stability of cortical networks and in brain dynamics [gerstner2002spiking, deco2014local].
The sparsity of representations has computational benefits for sensory processing. [olshausen2004sparse] underline that sparsity may be the optimal encoding strategy for neuronal networks, for it is energy efficient. Sparsity also increases the memory-storage capacity, and eases readout at subsequent processing layers. [babadi2014sparseness] show that sparse and expansive coding (i.e., from a lower dimensional sensory input space to a higher dimensional neural representation) reduces the intra-stimulus variability, maximises the inter-stimulus variability, and allows optimal and efficient readout of downstream neurons. This is the reason why sparse and expansive transformations are widespread in biology, e.g., in rodents [mombaerts1996visualizing] or flies [turner2008olfactory].
Overall, our findings – focused on the emerging properties of representations – corroborate the idea that Forward-Forward might be a better model than Backprop for learning in the cortex [hinton2022forward]. As shown in section 4, besides being a non biological learning rule, Backprop elicits non-sparse and less biologically plausible representations.
This work is a starting point for further explorations in the field of biologically plausible representations. Promising avenues for future research in this field encompass exploring model compression through pruning [blalock2020state], with the design of new strategies based on the relevance of the ensembles, as well as delving into the dynamic evolution of ensemble size and organisation throughout the training process. Such investigations hold the potential to shed light on the formation, evolving interactions, and persistence or replacement of ensembles. Furthermore, these findings can be cross-referenced with neurophysiological data to assess the biological plausibility of ensembles in dynamic learning contexts.
6 Limitations
The limitations of the present work could be addressed by applying similar analyses to more complex datasets and a variety of tasks. To scale to a more challenging dataset may require the replacement of a fully connected network with more suitable architectures trainable with the Forward-Forward protocol, e.g. the CNNs recently introduced in [Papachristodoulou_2024]. This approach could provide insights into how data characteristics influence sparsity levels and the resulting ensemble-like structures. Moreover, as highlighted in [yang2023theory], the choice of hyperparameters potentially affects representation sparsity, as well as the goodness function of choice. Future work should explore the effects of these factors to further address the limitations identified.
@@ 6 Limitations
Appendix A Computational resources
Appendix B Data
Svhn contains colored images of digits from house numbers, captured by Google StreetView. The images are composed of $32\times 32$ RGB-encoded pixels. This dataset is slightly larger than the previous two, as it contains $73257$ data-points in the training set and $26032$ in the test set.
The Svhn images have been cropped in order to center the digit of interest within the frame. However, the presence of adjacent digits and other distracting elements, that have been kept within the images, introduces an additional layer of complexity when compared to Mnist and FashionMnist, where the subjects are prominently displayed against a uniform black background. The CIFAR10 consists of $60000$ colored natural images categorized in $10$ balanced classes. The dataset is split in $50000$ training images and $10000$ test images. Each image, like Svhn has a resolution of $32\times 32$ for each channel. Compared to previous datasets, this is the most challenging one for a fully connected network. The dataset split employed is provided by the TorchVision framework [torchvision2016].
Appendix C Training details
All our models (FF, BP/FF and BP), on all datasets (Mnist, FashionMnist, Svhn and Cifar10), have been optimised using Adam [kingma2017adam] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ implemented in PyTorch [paszke2019pytorch]. A hyperparameter search has been performed to achieve sufficient accuracy for each model across all datasets. Every model was trained using batches of size $1024$.
Table 4: Hyperparameters selected to train our models.
Model Mnist FashionMnist Svhn CIFAR10
Epochs $1200$ $100$ $1000$ $1000$
FF Learning rate $0.01$ $0.01$ $0.0001$ $0.0001$
Epochs $300$ $300$ $200$ $200$
BP/FF Learning rate $0.0001$ $0.0001$ $0.0001$ $0.0001$
Epochs $80$ $80$ $80$ $80$
BP Learning rate $0.0001$ $0.0001$ $0.0001$ $0.0001$
Appendix D Activation patterns in deeper layers
In subsection 4.2 we claimed that in FF and BP/FF the images of a given category activate consistently a small set of units that we named ensembles, that share similarities to what is observed in sensory cortices. We reported in Figure 1 the activation map for Layer 1 (the first hidden layer) of FF trained on the Mnist dataset, and observed that very sparse ensembles emerge. In this section we show, in a similar fashion, the representations for Layers 2 and 3 (Figure 5 and Figure 6, respectively). We found that high sparsity also for deeper layers of this specific network; a qualitatively similar conclusion is reached for FF models trained on FashionMnist, Svhn and Cifar10. In BP/FF models a similar sparsity levels are observed, with the exception of the last layer that turns out to be not-sparse in each of the datasets we considered (see Table 3).
Figure 5: Activation patterns in a Multi-Layer Perceptron trained with the Forward-Forward algorithm, on the Mnist dataset. The image represents the activation map for neurons in Layer 2 for all images, grouped by class. A blue dot in position $(x,y)$ indicates that neuron $x$ is activated by input $y$; colorscale represents the intensity of such activation (incorrectly classified samples have been removed). Horizontal bands mark different categories; dark blue vertical lines mark active neurons. Each input category activates consistently a specific sets of neurons (ensemble). The sparsity measured according with the definition provided in subsection 3.5 is $0.84$.
Appendix E Activation patterns in different models
In Figure 1 Panel C we show the activation patterns in Layer 1 of FF trained on Mnist. For the purpose of a qualitative comparison, we show here analogous patterns for BP/FF and BP (see Figure 7 and Figure 8).
Appendix F Further results on representations of unseen categories and their ensembles
We showed in subsection 4.4 that a FF model trained on the FashionMnist dataset – deprived of one category – can respond at test time to this unseen category with a valid ensemble (Figure 4).
We report here the results of similar experiments, removing one category at a time. It turns out that, in each of the ten possible cases (we performed a single run for each category), the representations of the unseen category form a valid ensemble; we show three examples in Figure 9, different from the example shown in (Figure 4). It is with this situation in mind that we refer to “the ensembles related to unseen categories”.
When an unseen category forms a valid ensemble, it generally exhibits a high level of integration with the ensembles associated with the categories encountered during training. This integration implies that it can share common units with ensembles belonging to related categories. We show in Figure 10 how the ensembles of missing categories (same examples as in Figure 9) integrate – by sharing units – with the other ensembles.
Overall, these result relates to biological neural networks [yoshida2020natural, yuste2015neuron], where ensembles appear to be the functional building block of brain representations even in the absence of known stimuli.
Appendix G Distribution of excitatory and inhibitory connections
As we observed in subsection 4.2 FF and BP/FF have comparable sparsity levels and –when they are defined, the ensembles have comparable sizes. The presence of sparse ensembles suggests that a strong inhibition mechanism is at work, leaving only a few neurons active for each data sample. Inhibition in these architectures is the result of an interplay between the sign/magnitude of the weights and the sign/magnitude of the biases. Therefore it is natural to ask the following question: “are the FF and BP/FF similar also in this interplay between weights and biases?”. We find that the answer is no: the FF and BP/FF are indeed two profoundly different models that create sparse representations and ensembles with different mechanisms.
To show this, we consider for each neuron $i$ in a layer with width $n$ the fraction of its positive weights w.r.t. the total number of its input connections ($\varrho_{i}^{+}$ in the following), and its bias $\beta_{i}$. From these neuron-level quantities we construct their layer averages: $\varrho^{+}=\frac{1}{n}\sum_{i}\varrho_{i}^{+}$ and $\beta=\frac{1}{n}\sum_{i}\beta_{i}$. The neuron’s weights are strongly imbalanced towards inhibition when $\varrho_{i}^{+}\approx 0$ and, viceversa, strongly imbalanced towards excitation when $\varrho_{i}^{+}\approx 1$; when its $\varrho_{i}^{+}\approx 0.5$ we will say that the neuron’s weights are almost perfectly balanced; similar considerations holds for the biases, where a large and negative $\beta_{i}$ means strong inhibition for the $i$-th unit.
Focusing on the second hidden layer, we observe macroscopic differences of the empirical distribution of $\varrho_{i}^{+}$ among all the three models (see 11), with 1) a dominance of positive weights in the case of FF, 2) a bimodal distribution of $\varrho_{i}^{+}$ in BP/FF, with two populations of unbalanced units in opposite directions, and 3) a unimodal and approximately balanced distribution for all the neurons in the BP model.
[…]
Therefore, although the weights of the FF model appear imbalanced towards excitation, the average bias $\beta$ is very large and negative, and this might explain why, for this model, we observe such high sparsity values. On the other hand, the BP/FF model, have substantially zero average bias (within very small fluctuations), therefore the inhibitory mechanism at work here depends does not rely on a negative bias, but on the weights configuration.
From these results, we conclude that not only the training objective (i.e. goodness-based vs. categorical cross-entropy minimisation), but the specific training protocols (FF vs BP/FF) are determinant in creating different a interplay between excitation and inhibition, even when similar sparsity levels are observed.
@@ Appendix G Distribution of excitatory and inhibitory connections
The data shown in Figure 11 refer to the second hidden layer of models trained on Mnist. We report the statistics $\varrho^{+},\beta$ for all the combinations of models, layers and datasets in Table 5 and Table 6.
Table 5: Average fraction of positive weights ($\varrho^{+}$) for each combination of models, datasets and layers. Results are expressed as mean $\pm$ std. dev. over one single run.
@@
Table 6: Average bias ($\beta$) for each combination of models, datasets and layers. Results expressed as mean $\pm$ std. dev. over one single run.

### v4

@@ 4 Results
In this section, we describe our findings for the three models introduced, on the Mnist, FashionMnist, Svhn and Cifar10 datasets. In particular, we focus on the properties of representations obtained within the FF model, i.e., a model trained with Forward-Forward on its natural goodness objective. Such properties, such as the emergence of category-specific ensembles and the presence of shared units across them, establish a link between neural networks trained with the Forward-Forward algorithm and biological cortical networks described in subsection 2.2.
@@ 4.1 Classification accuracy
Before we present the main results of this work, we evaluate the performances of our models on the classification tasks at hand. Table 1 contains results in terms of test set classification accuracy for all models we employed – FF, BP/FF and BP – on Mnist, FashionMnist, Svhn and Cifar10. While some of these accuracy values are far from the state-of-the-art (i.e., respectively, $0.997$ (Cireşan et al., 2010), $0.931$ (Xiao et al., 2017), $0.860$ (Pitsios, 2017) and approximately $0.7$ (Lin et al., 2015), for fully-connected networks), they are a solid ground on which to build our subsequent investigations. Training details and hyperparameters for all models are reported in Appendix C.
@@ 4.2 Forward-Forward elicits sparse neuronal ensembles
Figure 2: Sparsity of category-specific representations. We report the sparsity of representations - computed as described in subsection 3.5 - for the three models FF, BP/FF and BP on the Mnist dataset. Sparsity values are the average over 10 runs.
The FF and BP/FF models – based on the original Forward-Forward network architecture, and trained according to the goodness objective (subsection 3.2 and subsection 3.3) – exhibit typically high sparsity levels in their representations, in clear contrast with BP (see Figure 2 and Table 2). While sparsity does not spontaneously arise in BP, it can be enforced by means of $\ell_{1}$ regularisation of the activations (Georgiadis, 2019). An analysis of this setting is presented in Appendix I.
Table 2: Average sparsity for all combinations of model, dataset and layer, according to the definition given in subsection 3.5. Results are expressed as mean $\pm$ std. dev. computed over 10 runs with independent random weights initialisation.
@@
When the sparsity level is sufficiently high ($S>0.5$) we are able to identify small sets of neurons (ensembles) that consistently co-activate across all the samples of the same class, similar to what has been observed in cortical representations (Yuste, 2015; Miller et al., 2014; Harris, 2005).
Figure 1 (Panels B, C) shows an example of average neuron activations for each class in Layer 1 of the FF model trained on Mnist, and showcases the emergence of sparse, category-specific, ensembles (see the Appendix E for a similar visualisation for Layers 2 and 3 of the same model and Appendix F for Layer 1 in all the models). These representations typically activate only a small fraction of units: ensembles consisting of just a few percent of the neurons in a layer are commonly observed, whether working with simpler datasets (e.g., Mnist, FashionMnist) or more complex ones (Svhn, Cifar10), with a slight tendency of the FF model to create larger ensembles in the latter case (Table 3).
@@
Overall, these findings show that networks trained with the Forward-Forward objective produce highly sparse representations, characterised by ensembles, i.e., small groups of neurons with category-specific activation patterns.
4.3 Visually similar classes can elicit ensembles with shared neurons
Drawing a parallel with a phenomenon observed in Neuroscience (Yoshida & Ohki, 2020), related categories can be expected to share units of their ensembles. This is indeed what we observe, as shown in Figure 3. Results are reported for FashionMnist, where different classes of clothes or shoes may contain a common share of visual features. In this regard, we observe a clear tendency to share units between similar classes – e.g., across representations of pullover, coat and shirt.
In the following section, we provide evidence that a unit can be shared across two ensembles even if one of these refers to an unseen category (i.e., excluded from the training set but whose representation, extracted at test time, generates an ensemble), as we show in Figure 4 (Panel C).
Figure 3: Visually similar classes in FashionMnist can elicit ensembles with shared neurons.
@@ 4.3 Visually similar classes can elicit ensembles with shared neurons
Panel B Element $i,j$ of the matrix indicates how many units are shared between the ensembles of category $i$ and category $j$ (normalised by the ensemble sizes), by using the Jaccard similarity index: $J(\mathcal{E}^{i},\mathcal{E}^{j})=\frac{\mid\mathcal{E}^{i}\cap\mathcal{E}^{j}\mid}{\mid\mathcal{E}^{i}\cup\mathcal{E}^{j}\mid}$. The results are referred to a single training run.
These findings indicate that ensembles relative to visually related classes can partially overlap.
@@ 4.4 Representations of unseen categories can elicit well-defined ensembles
We investigate the ability of a trained FF model to respond to unseen categories with a coherent activation pattern which is typical of the ensembles we found on the categories seen at training time. To this end, we repeatedly train FF on FashionMnist, removing one category at a time. Then, we extract the representation of the missing category, and verify if an ensemble is formed. We find that in all the ten cases, this is indeed the case, and the new ensemble share the same characteristics of the ones emerging for seen categories, apparently with the only exception of a lower average activation of their constituent units (see Figure 4 for one example, and the Appendix G for a more detailed account).
In several cases, we also find that the ensembles of unseen categories share units with the ensembles of seen categories, when endowed with similar visual features (Figure 4, Panel C). A more extensive exploration of these cases is also reported in the Appendix G.
These results show that ensemble-like structures can arise in a zero-shot setting on data categories held out of the training set.
Figure 4: The representations of an unseen category form an ensemble in FF trained on FashionMnist.
@@
4.5 Distribution of excitatory and inhibitory connections
As we observed in subsection 4.2, FF and BP/FF have comparable sparsity levels and, when they are defined, the ensembles have comparable sizes. A more fine-grained inspection of the representations learned by different models at different layers can be found in Appendix J. The presence of sparse ensembles suggests that a strong inhibition mechanism is at work, leaving only a few neurons active for each data sample. Inhibition in these architectures is the result of an interplay between the sign and magnitude of the weights and the those of the biases. Therefore it is natural to wonder whether FF and BP/FF are similar also in this interplay between weights and biases. We find that the answer is no: the FF and BP/FF are indeed two profoundly different models that create sparse representations and ensembles with different mechanisms.
To show this, we consider for each neuron $i$ in a layer with width $n$ the fraction of its positive weights w.r.t. the total number of its input connections ($\varrho_{i}^{+}$ in the following), and its bias $\beta_{i}$. From these neuron-level quantities we construct their layer averages: $\varrho^{+}=\frac{1}{n}\sum_{i}\varrho_{i}^{+}$ and $\beta=\frac{1}{n}\sum_{i}\beta_{i}$. The neuron’s weights are strongly imbalanced towards inhibition when $\varrho_{i}^{+}\approx 0$ and, viceversa, strongly imbalanced towards excitation when $\varrho_{i}^{+}\approx 1$; when $\varrho_{i}^{+}\approx 0.5$ we will say that the neuron’s weights are almost perfectly balanced; similar considerations hold for the biases, where a large and negative $\beta_{i}$ means strong inhibition for the $i$-th unit.
Focusing on the second hidden layer, we observe macroscopic differences in the empirical distribution of $\varrho_{i}^{+}$ among the three models (see Figure 5), with 1) a dominance of positive weights in the case of FF, 2) a bimodal distribution of $\varrho_{i}^{+}$ in BP/FF, with two populations of imbalanced units in opposite directions, and 3) a unimodal and approximately balanced distribution for all the neurons in the BP model.
@@ 4.5 Distribution of excitatory and inhibitory connections
We report the summary statistics $\varrho^{+},\beta$ for all the combinations of models, layers and datasets in Table 4 and Table 5. In all settings, we observe a similar picture, with the inhibition mechanisms dominated by negative biases in FF and negative weights in BP/FF.
In summary, these findings illustrate that although FF and BP/FF learn similar representations, they achieve sparsity by relying on fundamentally different neuron inhibition strategies.
Table 4: Average fraction of positive weights ($\varrho^{+}$) for each combination of models, datasets and layers. Results are expressed as mean $\pm$ std. dev. over a single training run.
@@
Table 5: Average bias ($\beta$) for each combination of models, datasets and layers. Results expressed as mean $\pm$ std. dev. over a single training run.